What is th numbe3r which when multiplie by 2 and added to 8 gives the same result as when it is divided by 2 and has 32 added to it?

Answer by ewatrrr(10682)


Hi,
Let x represent the number
Write as we Read***
2x + 8 = x/2 + 32 |Multiplying each term by 2 so as all denominators = 1
4x + 16 = x + 64
3x = 48
x = 16
CHECKING our Answer
32 + 8 = 8 + 32